OpenStudy (anonymous):

Help please, I'll love you foreverrrrr! xo "Paraffin, a wax used to make candles, has a molecular formula of C25H52 (s) + 38 O2 ---> 23 CO2 (g) + 26 H2O (l) How many liters of Co2 are produced when a 23.4g paraffin candle burns at STP?"

OpenStudy (aaronq):

Convert the mass given to moles (this requires adding up the molar masses of the individual elements - data given on the periodic table). the formula is: \(\sf moles=\dfrac{mass}{molar~mass}\) Next you need to use the molar ratio. \(\color{red}{1}~C_{25}H_{52} (s) + \color{red}{38}~ O_2 ---> \color{red}{23}~ CO_2 (g) + \color{red}{26}~ H_2O (l)\) Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients, \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) so for you equation it's: \(\dfrac{n_{CO_2}}{23}=\dfrac{n_{Parafin}}{1}\) plug in moles (the symbol n), solve for moles of \(CO_2\).

OpenStudy (aaronq):

After to convert the moles to Liters of gas, you need to use the ideal gas law, PV=nRT but for instances where you're at STP, there is a shortcut. Multiply the moles by 22.414 L, this is the volume of one mole of any gas at STP.